how do I graph 1/2x-3y+6
or what form is it in because then I may be able to figure it out myself

Answer by ewatrrr(24785)

 
Hi,
One can only graph an EQuation
If you meant...
1/2x-3y+6 = 0 ⇒ y=(1/6)x + 2 , Line m = 1/6, y-intercept = 2